OpenStudy (anonymous):

13 (X + 6) - 3 = 3X -5 10 7 -70 -8

OpenStudy (goformit100):

13 (X + 6) - 3 = 3X -5 or, 13X + 78 - 3 = 3X - 5 or, 13X - 3X = -78 - 5 or, 10X = -83 + 3 or, 10X = 80 OR, X = 80/10 = 8.............Ans.